Write an equation of the line that passes through each pair of points. Algebra 1
What is (-5/4 , 1) (-1/4 , 3/4)
Please help me find the slope and Y=
Also shows the step! :D
I would love your help.

Answer by nerdybill(7384)

(-5/4 , 1) (-1/4 , 3/4)
are two pairs of (x,y) points on a line:
(x1, y1) and (x2, y2)
.
Slope between two point is:
(y2-y1)/(x2-x1)
plug in the given values:
(3/4 - 1)/(-1/4 - (-5/4))
(3/4 - 4/4)/(-1/4 + 5/4)
(-1/4)/(4/4)
(-1)/(4)
-1/4 (slope) of line
.
Plug the newly found slope (-1/4) and one of the two given points (-5/4, 1) into the "point-slope" form of a line:
y - y1 = m(x - x1)
y - 1 = (-1/4)(x - (-5/4))
y - 1 = (-1/4)(x + 5/4)
y - 1 = (-1/4)x - 5/16
y = (-1/4)x - 5/16 + 1
y = (-1/4)x - 5/16 + 16/16
y = (-1/4)x + 11/16 (this is what they're looking for in "slope-intercept" form)